Could someone tell me if I understood this right? I should find out if these are geometric (if then I should give the common ratio) or arithmetic (if then I should give the common difference)
Are these right;
8,12,16,20... are Geometric with ratio=4
3,-6,12,-24...are geometric with ratio=-3
8,11,13,15... are none of both

Answer by Theo(2978)

Are these right;
-----
FIRST SEQUENCE
-----
8,12,16,20...
-----
this one is arithmetic with common difference = 4
12 - 8 = 4
16 - 12 = 4
20 - 16 = 4
8 + 4 = 12 + 4 = 16 + 4 = 20 ...
-----
SECOND SEQUENCE
-----
3,-6,12,-24...
-----
this one is geometric with common ratio = -2 (not 3)
-6 / 3 = -2
12 / -6 = -2
-24 / 12 = -2
3 * (-2) = -6 * (-2) = 12 * (-2) = -24 ...
-----
THIRD SEQUENCE
-----
8,11,13,15...
-----
can't see an arithmetic sequence here.
thought it would be 2 in an arithmetic but the first number throws it off since the difference between 11 and 8 is 3. once it gets past 11 it seem to be a common difference of 2 but the first number is an exception unless it was really 9?
11 - 8 = 3
13 - 11 = 2
15 - 13 = 2
as it is there doesn't seem to be a common difference
-----
can't see a geometric sequence here.
11/8 = 1.375
13/11 = 1.18181818...
15/13 = 1.153846...
there doesn't seem to be a common ratio
-----